40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 28 03:21:35 2010, Suresh, Update, Locking, Aligned the MC2 transmission photodiode 
    Reply  Thu Oct 28 04:23:38 2010, rana, Update, IOO, A2L prep 
Message ID: 3806     Entry time: Thu Oct 28 04:23:38 2010     In reply to: 3804
Author: rana 
Type: Update 
Category: IOO 
Subject: A2L prep 

To get the angle to length signal before the c1ioo processor gets going, we need a length signal. We can use either the error signal or the control signal.

I recommend using the control signal since its not puny. The 4-pin LEMO inputs to the OSEM ADC that Suresh has wired are differential so we can, in principle, use either the BNC output of the SERVO plug or the 2-pin LEMO output.

The analog whitening on the OSEM Whitening board should be engaged via the SUS MEDM screen so that we get a good SNR at the A2L dither frequencies.

If the ADC saturates, then we should use a pomona box RC low pass to cut everything off above 100 Hz.

Also, a comment about Yuta's elog: we estimated that the seismic motion was ~1e-7 - 1e-6 meters. The MC linewidth ought to be ~lambda/(2*Finesse) ~ 1e-9.

So, the MC servo as it was was not giving us enough gain (1/f above 50 Hz; UGF ~5-10 kHz) to get the error signal to stay in the linear PDH region. Kevin's filter gave us ~10x more gain at the seismic frequencies (1-3 Hz) of concern.

ELOG V3.1.3-